OpenStudy (anonymous):

is 4 4/9 divided by 7 1/5 equaled to 5/27 ?

hero (hero):

http://www.wolframalpha.com is your friend

OpenStudy (anonymous):

i think its 50/81

OpenStudy (anonymous):

is it (4)(4/9) or is it a mixed fraction?

OpenStudy (anonymous):

for mixed fractions its like this....

OpenStudy (anonymous):

\[4 \left(\begin{matrix}4 \\ 9\end{matrix}\right) divided by 7 \left(\begin{matrix}1 \\ 5\end{matrix}\right)\]

OpenStudy (anonymous):

then you have to find equivalent fractions for the 4 and the 7

OpenStudy (anonymous):

so multiply 9 by 4 and you get 36.

OpenStudy (anonymous):

multiply 5 by 7 and you get 35

OpenStudy (anonymous):

now the problem looks like this...

OpenStudy (anonymous):

\[\left[\begin{matrix}\left(\begin{matrix}35 \\ 5\end{matrix}\right) + \left(\begin{matrix}1 \\ 5\end{matrix}\right) \\ \left(\begin{matrix}36 \\ 9\end{matrix}\right) + \left(\begin{matrix}4 \\ 9\end{matrix}\right)\end{matrix}\right]\]

OpenStudy (anonymous):

add these up

OpenStudy (anonymous):

(36/5) / (40/9)

OpenStudy (anonymous):

then you take the inverse of (40/9) which is (9/40)

OpenStudy (anonymous):

actually its the other way around ...i messed up the division order

OpenStudy (anonymous):

so you multiply the inverse of 36/5 by 40/9 which is 200/324

OpenStudy (anonymous):

then simplify and you get 50/81

OpenStudy (anonymous):

if the 4 and 4/9 are multiplied its like this

OpenStudy (anonymous):

(4/1)(4/9) divided by (7/1)(1/5)

OpenStudy (anonymous):

so (12/9) divided by (7/5) so the answer is 20/21

OpenStudy (anonymous):

ok? hope this helped